Skip to main contentSkip to solution

Consider two independent events A and B such that P(A)=0.3P(A) = 0.3, P(B)=0.6P(B) = 0.6.

Match List-I with List-II

List-IList-II
(A) P(AP(A and B)B)(I) 0.28
(B) P(AP(A and not B)B)(II) 0.18
(C) P(AP(A or B)B)(III) 0.12
(D) PP(neither A nor B)(IV) 0.72

Choose the correct answer from the options given below:

Solution

✅ Correct Option: 3

Events A and B are independent with P(A)=0.3P(A) = 0.3 and P(B)=0.6P(B) = 0.6.

For independent events, P(A and B)=P(A)×P(B)P(A \text{ and } B) = P(A) \times P(B).


(A) P(A and B)=P(A)×P(B)P(A \text{ and } B) = P(A) \times P(B)

=0.3×0.6= 0.3 \times 0.6

=0.18= 0.18

(A) matches with (II)


(B) P(not B)=1−P(B)P(\text{not } B) = 1 - P(B)

=1−0.6= 1 - 0.6

=0.4= 0.4

P(A and not B)=P(A)×P(not B)P(A \text{ and not } B) = P(A) \times P(\text{not } B)

=0.3×0.4= 0.3 \times 0.4

=0.12= 0.12

(B) matches with (III)


(C) P(A or B)=P(A)+P(B)−P(A and B)P(A \text{ or } B) = P(A) + P(B) - P(A \text{ and } B)

=0.3+0.6−0.18= 0.3 + 0.6 - 0.18

=0.72= 0.72

(C) matches with (IV)


(D) P(not A)=1−P(A)P(\text{not } A) = 1 - P(A)

=1−0.3= 1 - 0.3

=0.7= 0.7

P(neither A nor B)=P(not A)×P(not B)P(\text{neither A nor B}) = P(\text{not } A) \times P(\text{not } B)

=0.7×0.4= 0.7 \times 0.4

=0.28= 0.28

(D) matches with (I)


Final matching:

  • (A) → (II) → 0.18
  • (B) → (III) → 0.12
  • (C) → (IV) → 0.72
  • (D) → (I) → 0.28

Keyboard Shortcuts

  • Left arrow: Previous question
  • Right arrow: Next question
  • S key: Jump to solution
  • Q key: Jump to question